OpenStudy (anonymous):

write in log form. 27=3^3

OpenStudy (agent0smith):

\[\Large a^x = b \text{ ...then... } \log_a b = x\]

OpenStudy (anonymous):

okay i got it.. then how do i find the value of log base 2 (16)

OpenStudy (agent0smith):

16 is the same as 2 to what power? 2^2 = 4 2^3 = 8 ...

OpenStudy (anonymous):

is that the answer?

OpenStudy (agent0smith):

Er, no...

OpenStudy (anonymous):

oh... what is it?

OpenStudy (agent0smith):

Try to answer the question I just asked.. 16 is the same as 2 to what power?

OpenStudy (anonymous):

4?

OpenStudy (agent0smith):

Yes, so log base 2 (16) is \[\Large \log_2 2^4\]

OpenStudy (agent0smith):

So do you know how to evaluate that?

OpenStudy (anonymous):

yeah i got it:) lol

OpenStudy (agent0smith):

Good :)

OpenStudy (anonymous):

a few more?

OpenStudy (agent0smith):

Post a new question for more cos I gotta eat anyway.

OpenStudy (anonymous):

ok